The perimeter of a rectangle is 50cm. if its length exceeds its breadth by 5cm,find the dimensions of the rectangle.
Answer by mananth(16946)

The perimeter of a rectangle is 50cm. if its length exceeds its breadth by 5cm,find the dimensions of the rectangle.
...
breadth =x
length = x+5
...
2*(L+B)= Perimeter
2*(x+x+5)=50
2*(2x+5)=50
/2
2x+5 = 25
-5 -5
2x=20
/2
x= 10 cm the width
length = 15 cm.
...
m.ananth@hotmail.ca
Happy New Year